OpenStudy (anonymous):

If the integral of f(x-c)dx = 11 from 3 to 5, then the integral of f(x)dx from 3-c to 5-c is? I will write the equation w/ teh equation box

OpenStudy (anonymous):

\[If \int\limits_{3}^{5} f(x-c) dx = 11\] then, \[\int\limits_{3-c}^{5-c} f(X) dx = ?\]

OpenStudy (zarkon):

do a simple substitution

OpenStudy (anonymous):

wat do u mean? wat shud i subsititute

OpenStudy (zarkon):

u=x-c

OpenStudy (anonymous):

so then du =1?

OpenStudy (zarkon):

du=dx

OpenStudy (anonymous):

then wat?

OpenStudy (zarkon):

have you done u-substitution before?

OpenStudy (anonymous):

yea but im still confused

OpenStudy (zarkon):

u(x)=x-c \[\int\limits_{3}^{5}f(x-c)dx=\int\limits_{u(3)}^{u(5)}f(u)du=\cdots\]

OpenStudy (anonymous):

is it 5-c - (3-c) = 2 - 2c

OpenStudy (anonymous):

can u help me continue teh problem??

OpenStudy (zarkon):

\[11=\int\limits_{3}^{5}f(x-c)dx=\int\limits_{u(3)}^{u(5)}f(u)du=\int\limits_{3-c}^{5-c}f(u)du\]

OpenStudy (anonymous):

so the expression integral of f(x) dx from 3-c to 5 - c is 11?

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

so there is no other number value for teh asnwer?

OpenStudy (anonymous):

its just to prove that the 2 expressions are equal

OpenStudy (zarkon):

they are equal

OpenStudy (anonymous):

but i dont understand i thought wen u integrate an expression u are supposed to get smthg.

OpenStudy (zarkon):

? ... you did ... it is 11

OpenStudy (anonymous):

oh ok i guess i was jsut confused about wat teh question was aking...sorry XD